# MAFA: A multi-agent framework for annotation

### v1

@@ 5.3 Evaluation metrics
  • NDCG@$k$: Normalized Discounted Cumulative Gain, which measures the ranking quality considering the positions of relevant FAQs.
@@ 6.1 Overall performance on bank dataset
Single-Agent (No Emb) 0.215 0.335 0.685 0.671 0.691 0.713
Single-Agent (Emb) 0.255 0.365 0.715 0.707 0.728 0.748
Single-Agent w/ Ans (No Emb) 0.220 0.350 0.690 0.712 0.703 0.713
Single-Agent w/ Ans (Emb) 0.270 0.380 0.730 0.722 0.743 0.763
MAFA (Standard) 0.320 0.410 0.855 0.765 0.782 0.798
MAFA (Specialized) 0.355 0.425 0.865 0.790 0.802 0.815
Our multi-agent framework with specialized few-shot examples significantly outperforms all baselines and individual agents across all metrics. Specifically, MAFA (Specialized) achieves a 23.5% improvement in Top-1 accuracy and a 41% improvement in MRR compared to the traditional BM25 approach. Even compared to the best individual agent (Single-Agent w/ Ans (Emb)), MAFA (Specialized) shows an 8.5% improvement in Top-1 accuracy and an 6.8% improvement in MRR.
@@ 6.3 Ablation study
To understand the contribution of each component, we conducted an ablation study by removing one component at a time. Table 6 shows the results on our bank dataset. The results show that each component contributes to the overall performance. The specialized few-shot examples contribute a 4.3% improvement in Top-1 accuracy, while the judge agent has the largest impact at 10.3%. This confirms the importance of both our specialized few-shot example strategy and the reranking step in our framework.
Table 6: Ablation Study (Top-1 Accuracy)
@@
Without Judge Agent 0.532
The ablation results on FiQA show a similar pattern, with the judge agent contributing the most to performance (8.0%), followed by specialized examples (2.8%). This consistency across datasets reinforces the robustness of our approach.
@@ 7 Conclusion and future work
In this paper, we presented a Multi-Agent Framework for FAQ annotation (MAFA) that combines multiple specialized agents with a judge agent to produce optimal FAQ rankings. We introduced a novel specialized few-shot example strategy that tailors examples to each agent’s strengths, significantly improving performance. Our experiments on a real-world bank dataset and public benchmarks (LCQMC and FiQA) show that this approach outperforms traditional methods and individual agents.
@@
Acknowledgments and Disclosure of Funding
We thank JP Morgan Chase for funding and supporting this research and providing the datasets used in our experiments.
Appendix A Appendix / supplemental material
@@ A.1 Judge agent design
The judge agent is a critical component of our framework, responsible for producing the final ranking of FAQs. In this section, we detail the design and implementation of this agent.
A.1.1 Prompt engineering
The prompt for the judge agent is carefully designed to elicit thoughtful analysis and accurate reranking. It includes:
1system_message = """You are an expert judge of FAQ relevance for our bank.
2Your task is to carefully analyze user utterances and determine which FAQs best address their needs.
3Be precise and thorough in your analysis, as banking customers depend on accurate information.
4Always consider both the semantic similarity and the practical relevance of each FAQ to the user’s query.
5When a user is asking about financial products, security features, or account management, prioritize exact matches.
6You must return your rankings in proper JSON format with detailed reasoning for each decision."""
7
8instructions = """
9You are tasked with reranking candidate FAQs based on their relevance to a user utterance for our bank.
10
11Given a user utterance and a list of candidate FAQs (with their original relevance scores), please rerank them based on your expert judgment. Consider:
121. Semantic similarity to the user’s query
132. Intent matching (what the user is trying to accomplish)
143. Specificity (how directly the FAQ addresses the user’s needs)
154. Banking domain knowledge (what would be most helpful for a banking customer)
16
17Provide your reasoning for each FAQ and assign a new relevance score from 0-100 for each.
18
19Your response must be in JSON format:
20{
21 "reranked_faqs": [
22 {
23 "faq": "FAQ Title",
24 "relevance_score": 95,
25 "reasoning": "Your reasoning for this ranking"
26 },
27 ...
28 ]
29}
30
31Return exactly 5 FAQs, ranked by relevance to the user’s query.
32"""
Listing 1: Judge Agent Prompt Structure
The prompt also includes the user utterance, candidate FAQs with their scores, agent predictions, examples from the training set, and FAQ descriptions.
A.1.2 Reasoning process
The judge agent is designed to follow a structured reasoning process:
@@ A.1.2 Reasoning process
    Intent analysis: Analyze the user’s utterance to identify their core intent and any implied needs.
@@
    Candidate assessment: Evaluate each candidate FAQ based on how well it addresses the identified intent.
@@
    Agent consensus consideration: Consider which FAQs were recommended by multiple agents and analyze why.
@@
    Answer content analysis: Review the answer content of each FAQ to determine if it provides the information the user needs.
@@
    Banking context application: Apply domain knowledge about banking to prioritize FAQs that address security, compliance, and customer service needs.
  6. 6.
    Final ranking determination: Synthesize all findings to produce a final ranked list with relevance scores.
A.2 Single agent prompts
In this section, we provide examples of the prompts used for our single agent implementations, illustrating how we implemented the ARQ-inspired structured reasoning approach described in Section 4.2.1.
A.2.1 FAQ mapper without answers
Both FAQ Mapper without answer context agents use the following prompt structure:
1system_message = """
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3IMPORTANT GUIDELINES:
41. Analyze the user’s intent thoroughly
52. Match the intent to the most relevant FAQs
63. Rank FAQs by relevance (0-100 scale)
74. Provide clear reasoning for each match
85. Return exactly 5 FAQs unless there are fewer relevant ones
96. Be precise - banking customers need accurate information
10"""
11user_message = """
12You will be given a user utterance and a list of available FAQs. Your task is to:
131. Analyze what the user is truly asking about (identify the core intent)
142. Search through the available FAQs for relevant matches
153. Rank the top 5 most relevant FAQs based on:
16 - Semantic similarity to the user’s intent
17 - Specificity to the question
18 - Likelihood of containing the information the user needs
194. Provide a confidence score (0-100) for each match
205. Explain your reasoning process
21For banking-related queries, consider:
22- Security concerns take priority
@@ 96. Be precise - banking customers need accurate information
23- Account access questions require specific authentication-related FAQs
24- Transaction questions should match to relevant transaction FAQs
25- General inquiries should match to general information FAQs
@@
26You must produce your analysis as a JSON object according to the following schema:
27{
28"user_utterance": "The original user utterance",
29"intent_analysis": "A thorough analysis of what the user is asking for, including likely underlying needs",
30"primary_banking_category": "The main banking category this query falls under (Account Management, Security, Transactions, etc.)",
31"relevant_faqs": [
32{
33"faq": "The title of the FAQ",
34"relevance_score": 95,
35"reasoning": "Detailed explanation of why this FAQ is relevant to the user’s query"
36},
37...
38],
39"confidence_in_mapping": "HIGH/MEDIUM/LOW",
40"explanation_of_confidence": "Brief explanation of why you’re confident or uncertain about these mappings",
41"recommended_clarification_question": "If confidence is MEDIUM or LOW, provide a question that would help clarify the user’s intent"
42}
@@
43If there are fewer than 5 relevant FAQs, only include those that are actually relevant.
44Available FAQs:
45[List of FAQs...]
46User Utterance: "{user_utterance}"
47"""’
Listing 2: FAQ Mapper Without Embeddings Prompt
A.2.2 FAQ mapper with answers
@@
Both the FAQ mapper with answers agents follow a similar structure but includes answer content:
1system_message = """
@@
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3IMPORTANT GUIDELINES:
41. Analyze the user’s intent thoroughly
@@
52. Match the intent to the most relevant FAQs based on both the FAQ question and its answer content
63. Rank FAQs by relevance (0-100 scale)
@@
74. Provide clear reasoning for each match, considering the full context of the FAQ including its answer
@@
85. Return exactly 5 FAQs unless there are fewer relevant ones
96. Be precise - banking customers need accurate information
10"""
11user_message = """
@@ 96. Be precise - banking customers need accurate information
12You will be given a user utterance and a list of available FAQs with their answers. Your task is to:
@@
131. Analyze what the user is truly asking about (identify the core intent)
142. Search through the available FAQs for relevant matches
153. Rank the top 5 most relevant FAQs based on:
16 - Semantic similarity to the user’s intent
17 - Specificity to the question
18 - Whether the answer content directly addresses the user’s needs
19 - Likelihood of containing the information the user needs
204. Provide a confidence score (0-100) for each match
215. Explain your reasoning process
22For banking-related queries, consider:
23- Security concerns take priority
@@
24- Account access questions require specific authentication-related FAQs
25- Transaction questions should match to relevant transaction FAQs
26- General inquiries should match to general information FAQs
@@
27You must produce your analysis as a JSON object according to the following schema:
28{
29"user_utterance": "The original user utterance",
30"intent_analysis": "A thorough analysis of what the user is asking for, including likely underlying needs",
31"primary_banking_category": "The main banking category this query falls under (Account Management, Security, Transactions, etc.)",
32"relevant_faqs": [
33{
34"faq": "The title of the FAQ",
35"relevance_score": 95,
36"reasoning": "Detailed explanation of why this FAQ is relevant to the user’s query, including how the answer content addresses the query"
37},
38...
39],
40"confidence_in_mapping": "HIGH/MEDIUM/LOW",
41"explanation_of_confidence": "Brief explanation of why you’re confident or uncertain about these mappings",
42"recommended_clarification_question": "If confidence is MEDIUM or LOW, provide a question that would help clarify the user’s intent"
43}
44Available FAQs with their Answers:
45[List of FAQs with answers...]
46User Utterance: "{user_utterance}"
47"""’
Listing 3: FAQ Mapper With Answers Prompt
A.2.3 ARQ vs. Chain-of-Thought implementation
@@
For comparison, our CoT implementation uses a similar prompt but without the structured JSON format:
1system_message = """
@@
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3"""
4user_message = """
5You will be given a user utterance and a list of available FAQs. Your task is to find the most relevant FAQs for the user’s query.
6First, think step-by-step about what the user is asking for. Analyze their intent carefully.
7Then, identify which FAQs in the provided list would best address their query.
8Rank the top 5 most relevant FAQs, providing your reasoning for each selection.
9Available FAQs:
10[List of FAQs...]
11User Utterance: "{user_utterance}"
12"""
Listing 4: Chain-of-Thought Prompt Example
@@
A.3 Error analysis
We analyzed cases where our framework failed to rank the correct FAQ first. The most common error types were:
  • Ambiguous queries: Queries that could be interpreted in multiple ways, e.g., "fash" (fashion vs. phishing).
  • Technical terminology: Queries using technical banking terms that were not well-represented in the training data.
These findings suggest areas for future improvement, such as better handling of ambiguous queries.
A.4 Monitoring and evaluation
Continuous monitoring and evaluation are essential for maintaining system performance. We recommend:
  • Implementing a feedback loop where user interactions are logged and analyzed
  • Regularly retraining the system with new examples
  • Conducting periodic human evaluations of system outputs
  • Setting up automated monitoring for anomalies or performance degradation
A.5 Public dataset details
A.5.1 LCQMC dataset
The Large-scale Chinese Question Matching Corpus (LCQMC) Liu et al. [2018] is designed for the task of question matching, which aims to determine whether two questions have the same intent. The dataset contains over 260,000 question pairs collected from Baidu Knows, a popular Chinese community question-answering service.
For our experiments, we adapted this dataset to our FAQ annotation task by:
  • Treating one question in each pair as a "user utterance"
  • Treating the other question as an "FAQ question"
  • Using the provided binary label (0 or 1) to determine relevance
  • Creating a pool of potential FAQ questions from the dataset
This adaptation allowed us to evaluate our framework on a large-scale question matching task in a different language, demonstrating its cross-lingual capabilities.
A.5.2 FiQA dataset
The FiQA dataset is a financial domain question answering dataset created for the Financial Question Answering Task of the WWW’18 conference. It contains user questions about financial topics along with expert-provided answers.
For our experiments, we adapted this dataset to our FAQ annotation task by:
  • Using the original questions as FAQ questions
  • Creating variations of these questions as user utterances
  • Ensuring that the variations maintain the same intent as the original questions
  • Using the associated answers as FAQ answers
This adaptation allowed us to evaluate our framework on a finance-specific dataset that is different from our internal banking dataset, providing insights into the framework’s domain generalization capabilities.
A.6 Judge agent analysis
We analyzed the judge agent’s reasoning to understand how it makes decisions. Table 8 shows examples of reasonings provided by the judge agent.
@@ A.6 Judge agent analysis
How much money do i have What is account preview? This FAQ directly explains how users can view limited account information, such as balances, before signing into the app, which is exactly what the user is asking about.
@@
A.7 Practical considerations
Deploying a multi-agent FAQ annotation system in a production banking environment requires careful consideration of several practical factors:
A.7.1 Latency and resource requirements
Our framework involves multiple LLM calls, which can introduce latency. Table 9 shows the average latency for different configurations.
Table 9: Latency Analysis
Configuration Avg. Latency (ms)
Single Agent (Direct) 450
Single Agent (w/ Embedding) 550
MAFA (Sequential) 2,800
MAFA (Parallel) 650
By parallelizing agent calls and implementing efficient caching strategies, we can reduce latency to acceptable levels for production use. Our parallel implementation reduces the overall latency from 2,800 ms to 650 ms, making it comparable to a single agent approach while providing significantly better accuracy.
A.7.2 Adapting to new domains
Our experiments on public benchmarks demonstrate that the MAFA framework can be adapted to new domains with relatively minimal effort. Key considerations for domain adaptation include:
  • Few-shot example selection: Selecting appropriate few-shot examples for the new domain is crucial for performance.
  • Domain-specific instructions: Updating the agent prompts to include domain-specific terminology and considerations.
  • Answer integration: For domains where context from answers is important, ensuring that this information is properly integrated into the system.
Our findings suggest that the core architecture of MAFA is domain-agnostic, with the specialized few-shot example strategy providing a mechanism for efficient adaptation to new domains.
A.8 Societal Impacts

### v3

@@ 5.3 Evaluation metrics
  • NDCG@$k$: Normalized Discounted Cumulative Gain, which measures the ranking quality considering the positions of FAQs.
@@ 6.1 Overall performance on bank dataset
Single-Agent (No Emb) 0.215 0.365 0.685 0.671 0.691 0.713
Single-Agent (Emb) 0.255 0.395 0.715 0.707 0.728 0.748
Single-Agent w/ Ans (No Emb) 0.220 0.450 0.690 0.712 0.703 0.713
Single-Agent w/ Ans (Emb) 0.270 0.480 0.730 0.722 0.743 0.763
MAFA (Standard) 0.320 0.610 0.855 0.765 0.782 0.798
MAFA (Specialized) 0.355 0.625 0.865 0.790 0.802 0.815
Our multi-agent framework with specialized few-shot examples significantly outperforms all baselines and individual agents across all metrics. Specifically, MAFA (Specialized) achieves a 23.5% improvement in Top-1 accuracy and a 41% improvement in MRR compared to the traditional BM25 approach. Even compared to the best individual agent (Single-Agent w/ Ans (Emb)), MAFA (Specialized) shows improvements of 8.5% in Top-1 accuracy and an 6.8% in MRR.
@@ 6.2 Performance on public benchmarks
6.3 Fine-tuning considerations
While our framework achieves strong performance without fine-tuning, we recognize the potential benefits of domain-specific adaptation. Fine-tuning the embedding models on banking-specific question pairs could further improve retrieval accuracy, particularly for technical terminology. Similarly, fine-tuning the LLMs on banking FAQ data could enhance their understanding of domain-specific concepts and improve reasoning quality.
However, we chose not to pursue fine-tuning in this work for several reasons: (1) to demonstrate the framework’s effectiveness with off-the-shelf models, making it more accessible for practitioners; (2) to avoid overfitting to specific banking terminology that may vary across institutions; and (3) to maintain the generalization capabilities demonstrated on public benchmarks. Future work could explore selective fine-tuning strategies that preserve generalization while improving domain-specific performance.
@@ 6.4 Ablation study
To understand the contribution of each component, we conducted an ablation study by removing one component at a time. Table 6 shows the results on our bank dataset. The results show that each component contributes to the overall performance. The specialized few-shot examples contribute a 3.5% improvement in Top-1 accuracy, while the judge agent has the largest impact at 8%. This confirms the importance of both our specialized few-shot example strategy and the reranking step in our framework.
Table 6: Ablation Study on Bank Dataset (Top-1 Accuracy)
@@
Without Judge Agent 0.545
The ablation results on FiQA show a similar pattern, with the judge agent contributing the most to performance (6.7%), followed by specialized (diverse) few-shot examples (2.8%). This consistency across datasets reinforces the robustness of our approach.
@@ 7 Conclusion and future work
In this paper, we presented a Multi-Agent Framework for FAQ annotation (MAFA) that combines multiple ranker agents with a judge agent to produce optimal FAQ rankings for user utterances. We introduced a novel specialized few-shot example strategy that tailors examples to each agent’s strengths, significantly improving performance. Our experiments on a real-world bank dataset and public benchmarks (LCQMC and FiQA) show that this approach outperforms traditional methods and individual agents.
@@
{ack}
We would like to thank the Self Service Enablement team at JP Morgan Chase for their support and guidance throughout this research.
Appendix A Additional Implementation Details
@@ A.1 Judge Agent Design
The judge agent plays a pivotal role in the final ranking of FAQs, ensuring that users receive the most relevant information. Its prompt is meticulously crafted to encourage a comprehensive analysis: "You are an expert judge of FAQ relevance for our bank. Your task is to carefully analyze user utterances and determine which FAQs best address their needs. Be precise and thorough in your analysis, as banking customers depend on accurate information. Always consider both the semantic similarity and the practical relevance of each FAQ to the user’s query. When a user is asking about financial products, security features, or account management, prioritize exact matches. You must return your rankings in proper JSON format with detailed reasoning for each decision…" The judge agent adheres to a structured reasoning process, which includes several key steps. First, it conducts an Intent Analysis to identify the user’s core intent and implied needs. Next, it performs a Candidate Assessment to evaluate how well each candidate addresses the identified intent. The agent also considers Agent Consensus, analyzing FAQs recommended by multiple agents. It then conducts an Answer Content Analysis to assess the relevance of FAQ answers. Additionally, the agent applies Banking Context to prioritize FAQs using domain knowledge. Finally, it synthesizes these findings to produce a Final Ranking, ensuring that the most pertinent FAQs are presented to the user.
A.2 Public Dataset Details
A.2.1 LCQMC Dataset
The Large-scale Chinese Question Matching Corpus is designed for question matching, containing over 260,000 question pairs. We adapted this dataset by treating one question as a "user utterance" and the other as an "FAQ question," using the provided binary label to determine relevance.
A.2.2 FiQA Dataset
The FiQA dataset contains financial domain questions with expert-provided answers. We adapted this dataset by using original questions as FAQ questions and creating variations as user utterances.
A.3 Latency and Resource Requirements
Our framework involves multiple LLM calls, which can introduce latency. By parallelizing agent calls and implementing efficient caching, we reduced the overall latency from 2,800ms (sequential) to 650ms (parallel), making it comparable to single-agent approaches while providing significantly better accuracy.
A.4 Error Analysis
In our error analysis, we examined instances where our framework did not successfully rank the correct FAQ as the top result. The most prevalent types of errors were ambiguous queries and technical terminology. Ambiguous queries are those that can be interpreted in multiple ways, such as "fash," which could refer to either fashion or phishing. Technical terminology errors occur when queries use specialized banking terms that are not adequately represented in the training data. These findings highlight potential areas for future improvement, including enhancing the framework’s ability to handle ambiguous queries more effectively.
A.5 Single Agent Prompts
In this section, we present examples of the prompts used for our single agent implementations, demonstrating the ARQ-inspired structured reasoning approach described in Section 4.2.1.
A.5.1 FAQ Mapper Without Answers
Both FAQ Mapper agents without answer context utilize the following prompt structure:
System Message: You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
  • Analyze the user’s intent thoroughly.
  • Match the intent to the most relevant FAQs.
  • Rank FAQs by relevance (0-100 scale).
  • Provide clear reasoning for each match.
  • Return exactly 5 FAQs unless there are fewer relevant ones.
  • Be precise - banking customers need accurate information.
User Message: You will be given a user utterance and a list of available FAQs. Your task is to:
@@ A.5.1 FAQ Mapper Without Answers
    Analyze what the user is truly asking about (identify the core intent).
@@
    Search through the available FAQs for relevant matches.
@@
    Rank the top 5 most relevant FAQs based on:
    • Semantic similarity to the user’s intent.
    • Specificity to the question.
    • Likelihood of containing the information the user needs.
@@
    Provide a confidence score (0-100) for each match.
@@
    Explain your reasoning process.
For banking-related queries, consider:
  • Security concerns take priority.
@@
  • Account access questions require specific authentication-related FAQs.
  • Transaction questions should match to relevant transaction FAQs.
  • General inquiries should match to general information FAQs.
@@
You must produce your analysis as a JSON object according to the following schema:
"user_utterance": "The user utterance",
"intent_analysis": "A thorough analysis..",
"primary_banking_category": "..",
"relevant_faqs": [
"faq": "The title of the FAQ",
"relevance_score": 0-100,
"reasoning": "Detailed explanation"
"confidence_in_mapping": "HIGH/MEDIUM/LOW",
"explanation_of_confidence": "..",
"recommended_clarification_question": ".."
@@
If there are fewer than 5 relevant FAQs, only include those that are actually relevant.
A.5.2 FAQ Mapper With Answers
@@
Both FAQ Mapper agents with answers follow a similar structure but include answer content:
@@
System Message: You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
  • Analyze the user’s intent thoroughly.
@@
  • Match the intent to the most relevant FAQs based on both the FAQ question and its answer content.
  • Rank FAQs by relevance (0-100 scale).
@@
  • Provide clear reasoning for each match, considering the full context of the FAQ including its answer.
@@
  • Return exactly 5 FAQs unless there are fewer relevant ones.
  • Be precise - banking customers need accurate information.
@@
User Message: You will be given a user utterance and a list of available FAQs with their answers. Your task is to:
  1. 1.
@@
    Analyze what the user is truly asking about (identify the core intent).
  2. 2.
    Search through the available FAQs for relevant matches.
  3. 3.
    Rank the top 5 most relevant FAQs based on:
    • Semantic similarity to the user’s intent.
    • Specificity to the question.
    • Whether the answer content directly addresses the user’s needs.
    • Likelihood of containing the information the user needs.
  4. 4.
    Provide a confidence score (0-100) for each match.
  5. 5.
    Explain your reasoning process.
For banking-related queries, consider:
  • Security concerns take priority.
@@
  • Account access questions require specific authentication-related FAQs.
  • Transaction questions should match to relevant transaction FAQs.
  • General inquiries should match to general information FAQs.
@@
You must produce your analysis as a JSON object according to the following schema.. (same schema outlined in section A.5.1).
A.5.3 ARQ vs. Chain-of-Thought Implementation
@@
For comparison, our Chain-of-Thought (CoT) implementation uses a similar prompt but without the structured JSON format:
@@
System Message: You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
User Message: You will be given a user utterance and a list of available FAQs. Your task is to find the most relevant FAQs for the user’s query. First, think step-by-step about what the user is asking for. Analyze their intent carefully. Then, identify which FAQs in the provided list would best address their query. Rank the top 5 most relevant FAQs, providing your reasoning for each selection.
@@
A.6 Agent reasoning analysis
We analyzed the agents’ reasoning to understand how they make decisions. Table 8 shows examples of reasonings provided by the judge agent.
@@ A.6 Agent reasoning analysis
How much do i have What is account preview? This FAQ directly explains how users can view limited account information, such as balances, before signing into the app, which is exactly what the user is asking about.
@@
A.7 Societal impacts